The paper is asking for me to write the equation of the line whose slope is 1/5 and whose y-intercept is 4. How would I do this and show the work?
Answer by Alan3354(69443)

The paper is asking for me to write the equation of the line whose slope is 1/5 and whose y-intercept is 4. How would I do this and show the work?
----------
y = mx + b is the slope-intercept form
m is the slope, and b is the y-intercept
---------
The eqn is
y = (1/5)x + 4
Not much work to show.
--------
You can put it in standard form:
x - 5y = -20